# Higher Gaussian maps on the hyperelliptic locus and second fundamental form

### v1

@@ 6. Isotropic subspaces
By Lemma 4.2, with our choice of the basis $\omega_{1},...,\omega_{g-1}$ of $H^{0}(M)$, we have
@@
$\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=0\quad\forall h+l\leq 4k+1.$
@@
$\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=\sum\limits_{1\leq i<j\leq g-1,j\geq g-\tilde{l}}a_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=$
$=\sum\limits_{1\leq i<g-1}a_{ig-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)+\sum\limits_{1\leq i<g-2}a_{ig-2}(\omega_{i}^{(h)}\omega_{g-2}^{(l)}-\omega_{i}^{(l)}\omega_{g-2}^{(h)})(p)\\ +\dots+\sum\limits_{1\leq i<g-\tilde{l}}a_{ig-\tilde{l}}(\omega_{i}^{(h)}\omega_{g-\tilde{l}}^{(l)}-\omega_{i}^{(l)}\omega_{g-\tilde{l}}^{(h)})(p).$
Let us focus on the first term of the above sum. From Lemma 6.2, $Q\in Ker(\mu_{2k})$ implies $a_{ig-1}=0\quad\forall g-(2k+1)\leq i\leq g-2$. Then we have
(45) $\sum\limits_{1\leq i<g-1}a_{ig-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)=\sum\limits_{i=1}^{g-2k-2}a_{ig-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p).$
@@
Let us now consider the general term, where $j=g-s$. From Lemma 6.2, we have $a_{ig-s}=0$ $\forall g-(2k+2-s)\leq i\leq g-s$. Hence the $i$’s that actually appear in the sum are those satisfying $i\leq g-(2k+2-s)-1=g-2k-3+s$. In the same way as we did for the first term, for $h\leq 4k+3-2s$, we have $2i+3\leq 2(g-2k-3+s)\leq 2g-h$, so $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$ and the term is zero.
@@
Let $C$ be an hyperelliptic curve of genus $g$, $p\in C$ a Weierstrass point. Let $\gamma_{1},...,\gamma_{r}\in H^{0}(K_{C})$ where locally $\gamma_{\alpha}=g_{\alpha}dz_{\alpha}$ and let
$Q=\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta}=\sum\limits_{1\leq i<j\leq g-1}a_{ij}Q_{ij}\in Ker(\mu_{2k}).$
@@
$\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=0\quad\forall h+l\leq 4k+3.$
@@
(46) $\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=\\ =\sum\limits_{1\leq i<j\leq g-1}a_{ij}[(s\omega_{i})^{(h)}(t\omega_{j})^{(l)}+(t\omega_{j})^{(h)}(s\omega_{i})^{(l)}-(s\omega_{j})^{(h)}(t\omega_{i})^{(l)}-(t\omega_{i})^{(h)}(s\omega_{j})^{(l)}](p).$
@@
(47) $\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p).$
@@
$\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)=0\quad\forall m+n\leq 4k+1.$
@@
Let $Q=\sum\limits_{\alpha,\beta=1}^{r}b_{\alpha,\beta}\gamma_{\alpha}\otimes\gamma_{\beta}$, where $\gamma_{1},...,\gamma_{r}\in H^{0}(K_{C})$. We will prove that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$. By Remark 6.1, we know that
(48) $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=2\pi i\sum\limits_{u=0}^{2k}\left[\sum\limits_{\alpha,\beta=1}^{r}b_{\alpha,\beta}\gamma_{\alpha}^{(4k+4-u)}(0)\gamma_{\beta}^{(s)}(0)\right]\frac{(2k+1-u)}{u!(4k+4-u)!}$
@@
$\sum\limits_{\alpha,\beta=1}^{r}b_{\alpha,\beta}\gamma_{\alpha}^{(4k+4-u)}(0)\gamma_{\beta}^{(u)}(0).\quad(*)$
@@
$\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)$
@@
    $(st^{\prime\prime}-ts^{\prime\prime})(p)\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l-2)}-\omega_{j}^{(h)}\omega_{i}^{(l-2)})(p).$
@@
    $(s^{\prime\prime}t-t^{\prime\prime}s)(p)\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(h-2)}\omega_{j}^{(l)}-\omega_{j}^{(h-2)}\omega_{i}^{(l)})(p).$
@@
  • Assume $l$ is odd. We claim that we get zero. In fact, if $j\leq g-\frac{l+1}{2}$, then $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. Hence we can assume $j\geq g-\frac{l+1}{2}+1=g-\frac{l-1}{2}$. Let $j:=g-u$, hence $u\in\{1,...,\frac{l-1}{2}\}$. From Lemma 6.2 we have $a_{i,g-u}=0$ for $i\geq g-2k-2+u$, so we may assume $i\leq g-2k-3+u$.
@@
    We have that $\omega_{i}^{(h)}(p)=0$ if $i\leq g-\frac{h}{2}-\frac{3}{2}<g-\frac{h}{2}-1$, so $\omega_{i}^{(h)}(p)=0$ if $i\leq g-\frac{h}{2}-2$. Hence we can assume $i\geq g-\frac{h}{2}-1$, but $i\leq g-2k-3+u$ and $g-2k-3+u-(g-\frac{h}{2}-1)=u-2-2k+\frac{h}{2}\leq\frac{l}{2}+\frac{h}{2}-2-2k=0$ because $h+l=4k+4$. Hence if $u=l/2$, there is a unique nonzero term, which is precisely $a_{g-\frac{h}{2}-1,g-\frac{l}{2}}\omega_{g-\frac{h}{2}-1}^{(h)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)$.
@@
So let us assume $l$ even, then also $h$ is even. By Remark 6.3, $\omega_{j}^{(l-2)(p)}=0$ if $j\leq g-\frac{l}{2}-\frac{3}{2}<g-\frac{l}{2}-1$, so we may assume $j\geq g-\frac{l}{2}-1$. As usual we put $j=g-u$ and we have $u\leq\frac{l}{2}+1$. We have $\omega_{i}^{(l)}(p)=0$ if $32i\leq 2g-l-$ 3, but, as in the previuos case, we can assume $i\leq g-2k-3+u$. Since $l\leq 2k$, and $u\leq\frac{l}{2}+1$, we have $32g-6-4k+2u\leq 2g-l-$ 3. So $32i\leq 2g-6-4k+2u\leq 2g-l-$ 3, and hence $\omega_{i}^{(l)}(p)=0$.
Again, by Remark 6.3, $\omega_{i}^{(h-2)}(p)=0$ if $12i\leq 2g-h-$ 1. As usual we may assume $2i\leq 2g-6-4k+2u$. We have $u\leq\frac{l}{2}+1$. We claim that the only nonzero term occurs for $u=\frac{l}{2}+1$ and $i=g-\frac{h}{2}$. In fact, if $u\leq\frac{l}{2}$, we have $12g-6-4k+2u\leq 2g-h-$ 1, so $\omega_{i}^{(h-2)}(p)=0$. So $u=\frac{l}{2}+1$, $i=g-\frac{h}{2}$ and the only term is $a_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$
We have computed all the terms of (*). These are
  • $(st^{\prime\prime}-ts^{\prime\prime})(p)[\binom{l}{2}a_{g-\frac{h}{2}-1,g-\frac{l}{2}}\omega_{g-\frac{h}{2}-1}^{(h)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)-\binom{h}{2}a_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)]$ if $l\geq 2$ and $h,l$ even;
  • $-(st^{\prime\prime}-ts^{\prime\prime})(p)\binom{h}{2}a_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$ if $h$ is even and $l=0$.
@@
$\rho(Q)(\xi_{p}^{2k+3},\xi_{p}^{2k+1})=2\pi i\sum_{\begin{subarray}{c}2\leq l\leq 2k\\ l\text{ even}\end{subarray}}\frac{(2k+1-l)}{l!(4k+4-l)!}(st^{\prime\prime}-ts^{\prime\prime})(p)\cdot\\ \left[\binom{l}{2}a_{g-3-2k+\frac{l}{2},g-\frac{l}{2}}\omega_{g-3-2k+\frac{l}{2}}^{(4k+4-l)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)-\binom{h}{2}a_{g-2-2k+\frac{l}{2},g-\frac{l}{2}-1}\omega_{g-2-2k+\frac{l}{2}}^{(4k+2-l)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)\right]-\\ -2\pi i(st^{\prime\prime}-ts^{\prime\prime})(p)\frac{2k+1}{(4k+4)!}\binom{4k+4}{2}a_{g-2k-2,g-1}\omega_{g-2k-2}^{(4k+2)}(p)\omega_{g-1}(p)$
Notice that every term in the equation is a linear combination of $a_{r,u}$ where $r+u=2g-2k-3$.
Let $\lambda_{k,u}$ be the coefficient of $a_{g-2k-3+u,g-u}$ where $u=1,...,k+1$. We will now see that $\lambda_{k,u}\neq 0$ for every $k,u$. For $1\leq u\leq k$, we have
@@
(49) $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=\sum\limits_{u=1}^{k+1}\lambda_{k,u}a_{g-2k-3+u,g-u}=0$
for every $Q\in Ker(\mu_{2k})$. Recall that $Q\in Ker(\mu_{2k})$ if and only if the coefficients $a_{ij}$ satisfy the equations given in Lemma 3.7. Moreover observe that from Theorem 3.1 it follows that we have $k$ linearly independent equations in the variables $a_{g-2k-3+u,g-u}$ with $u=1,...,k+1$. We have shown that each coefficient $\lambda_{k,u}$ in equation (49) is different from zero and up to scalar it is the product $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)$. So we can multiply the sections $\omega_{g-u}$ by a non zero scalar multiple, hence equation (49) cannot be linearly dependent from the $k$ equations giving $Ker(\mu_{2k})$. Hence there exists a quadric $Q\in ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$.
@@
Notice that we have just proven that a quadric $Q\in Ker(\mu_{2k})$ satisfies $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=0,$ if and only if the coefficients of $Q$ satisfy the equations $a_{g-2k+3+u,g-u}=0,\forall u=1,...,k+1$ in addition to all the other equations in Lemma 3.7 defining $Ker(\mu_{2k})$. The condition $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=0$ gives a hyperplane in $Ker(\mu_{2k})$ that we denote by
@@
Let $0\leq k\leq\lfloor\frac{g-3}{2}\rfloor$ and let us now assume $Q\in A_{k,0}$. Then
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=\sum\limits_{u=1}^{k+1}\alpha_{k,u}a_{g-4-2k+u,g-u}\omega_{g-4-2k+u}^{(4k+6-2u)}(p)\omega_{g-u}^{(2u-2)}(p),$
where all the coefficients $\alpha_{k,u}$ are different from zero. Moreover,
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\Leftrightarrow a_{g-4-2k+u,g-u}=0\quad\forall u=1,...,k+1.$
@@
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=2\pi i[\sum\limits_{u=0}^{2k+2}\sum\limits_{\alpha,\beta=1}^{r}b_{\alpha,\beta}\gamma_{\alpha}^{(4k+6-u)}(0)\gamma_{\beta}^{(s)}(0)]\frac{(2k+3-u)}{u!(4k+6-u)!}.$
@@
(50) $\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)$
@@
$\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{i}^{(m)}\omega_{j}^{(n)})(p)=0\quad\forall m+n\leq 4k+3.$
@@
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=\sum\limits_{u=1}^{k+1}\alpha_{k,u}a_{g-4-2k+u,g-u}\omega_{g-4-2k+u}^{(4k+6-2u)}(p)\omega_{g-u}^{(2u-2)}(p),$
@@
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\Leftrightarrow a_{g-4-2k+u,g-u}=0\quad\forall u=1,...,k+1.$
@@
To show that these are the only ones, we will prove the result when $g$ is even, the case $g$ odd been very similar.

### v2

@@ 6. Isotropic subspaces
Let $\omega_{1},...,\omega_{g-1}$ be our basis of $H^{0}(M)$ and let $z$ be a local coordinate centered in the Weierstrass point $p$. Assume that we locally have $\omega_{i}=g_{i}(z)dz\otimes l^{-1}$, by abuse of notation we will write $\omega_{i}^{(h)}(p)$ instead of $g_{i}^{(h)}(0)$. Then by Remark 4.2 we have
@@
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=0\quad\forall h+l\leq 4k+1.$
@@
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=\sum\limits_{1\leq i<j\leq g-1,j\geq g-\tilde{l}}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=$
$=\sum\limits_{1\leq i<g-1}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)+\sum\limits_{1\leq i<g-2}b_{i,g-2}(\omega_{i}^{(h)}\omega_{g-2}^{(l)}-\omega_{i}^{(l)}\omega_{g-2}^{(h)})(p)\\ +\dots+\sum\limits_{1\leq i<g-\tilde{l}}b_{i,g-\tilde{l}}(\omega_{i}^{(h)}\omega_{g-\tilde{l}}^{(l)}-\omega_{i}^{(l)}\omega_{g-\tilde{l}}^{(h)})(p).$
Let us focus on the first term of the above sum. From Lemma 6.2, $Q\in Ker(\mu_{2k})$ implies $b_{i,g-1}=0\quad\forall g-(2k+1)\leq i\leq g-2$. Then we have
(50) $\sum\limits_{1\leq i<g-1}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)=\sum\limits_{i=1}^{g-2k-2}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p).$
@@
Let us now consider the general term, where $j=g-s$. From Lemma 6.2, we have $b_{i,g-s}=0$ $\forall g-(2k+2-s)\leq i<g-s$. Hence the $i$’s that actually appear in the sum are those satisfying $i\leq g-(2k+2-s)-1=g-2k-3+s$. In the same way as we did for the first term, for $h\leq 4k+3-2s$, we have $2i+3\leq 2(g-2k-3+s)\leq 2g-h$, so $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$ and the term is zero.
@@
Let $C$ be an hyperelliptic curve of genus $g$, $p\in C$ a Weierstrass point. Let $\gamma_{1},...,\gamma_{r}\in H^{0}(K_{C})$ where locally $\gamma_{\alpha}=g_{\alpha}dz$ and let
$Q=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta}=\sum\limits_{1\leq i<j\leq g-1}b_{ij}Q_{ij}\in Ker(\mu_{2k}).$
@@
$\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=0\quad\forall h+l\leq 4k+3.$
@@
(51) $\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=\\ =\sum\limits_{1\leq i<j\leq g-1}b_{ij}[(s\omega_{i})^{(h)}(t\omega_{j})^{(l)}+(t\omega_{j})^{(h)}(s\omega_{i})^{(l)}-(s\omega_{j})^{(h)}(t\omega_{i})^{(l)}-(t\omega_{i})^{(h)}(s\omega_{j})^{(l)}](p).$
@@
(52) $\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p).$
@@
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)=0\quad\forall m+n\leq 4k+1.$
@@
Let $Q=\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}\gamma_{\beta}$, where $\gamma_{1},...,\gamma_{r}\in H^{0}(K_{C})$. We will prove that there exists $Q$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$. By Remark 6.1, we know that
(53) $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=2\pi i\sum\limits_{u=0}^{2k}\left[\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}^{(4k+4-u)}(0)\gamma_{\beta}^{(u)}(0)\right]\frac{(2k+1-u)}{u!(4k+4-u)!}$
@@
(54) $\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}^{(4k+4-u)}(0)\gamma_{\beta}^{(u)}(0).\quad$
@@
$\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)$
@@
    $(st^{\prime\prime}-ts^{\prime\prime})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l-2)}-\omega_{j}^{(h)}\omega_{i}^{(l-2)})(p).$
@@
    $(s^{\prime\prime}t-t^{\prime\prime}s)(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h-2)}\omega_{j}^{(l)}-\omega_{j}^{(h-2)}\omega_{i}^{(l)})(p).$
@@
  • Assume $l$ is odd. We claim that we get zero. In fact, if $j\leq g-\frac{l+1}{2}$, then $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. Hence we can assume $j\geq g-\frac{l+1}{2}+1=g-\frac{l-1}{2}$. Let $j:=g-u$, hence $u\in\{1,...,\frac{l-1}{2}\}$. From Lemma 6.2 we have $b_{i,g-u}=0$ for $i\geq g-2k-2+u$, so we may assume $i\leq g-2k-3+u$.
@@
    We have that $\omega_{i}^{(h)}(p)=0$ if $i\leq g-\frac{h}{2}-\frac{3}{2}<g-\frac{h}{2}-1$, so $\omega_{i}^{(h)}(p)=0$ if $i\leq g-\frac{h}{2}-2$. Hence we can assume $i\geq g-\frac{h}{2}-1$, but $i\leq g-2k-3+u$ and $g-2k-3+u-(g-\frac{h}{2}-1)=u-2-2k+\frac{h}{2}\leq\frac{l}{2}+\frac{h}{2}-2-2k=0$ because $h+l=4k+4$. Hence if $u=l/2$, there is a unique nonzero term, which is precisely $b_{g-\frac{h}{2}-1,g-\frac{l}{2}}\omega_{g-\frac{h}{2}-1}^{(h)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)$.
@@
So let us assume $l$ even, then also $h$ is even. By Remark 6.3, $\omega_{j}^{(l-2)(p)}=0$ if $j\leq g-\frac{l}{2}-\frac{3}{2}<g-\frac{l}{2}-1$, so we may assume $j\geq g-\frac{l}{2}-1$. As usual we put $j=g-u$ and we have $u\leq\frac{l}{2}+1$. We have $\omega_{i}^{(l)}(p)=0$ if $32i\leq 2g-l-$ 3, but, as in the previous case, we can assume $i\leq g-2k-3+u$. Since $l\leq 2k$, and $u\leq\frac{l}{2}+1$, we have $32g-6-4k+2u\leq 2g-l-$ 3. So $32i\leq 2g-6-4k+2u\leq 2g-l-$ 3, and hence $\omega_{i}^{(l)}(p)=0$.
Again, by Remark 6.3, $\omega_{i}^{(h-2)}(p)=0$ if $12i\leq 2g-h-$ 1. As usual we may assume $2i\leq 2g-6-4k+2u$. We have $u\leq\frac{l}{2}+1$. We claim that the only nonzero term occurs for $u=\frac{l}{2}+1$ and $i=g-\frac{h}{2}$. In fact, if $u\leq\frac{l}{2}$, we have $12g-6-4k+2u\leq 2g-h-$ 1, so $\omega_{i}^{(h-2)}(p)=0$. So $u=\frac{l}{2}+1$, $i=g-\frac{h}{2}$ and the only term is $b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$
We have computed all the non-zero terms of (54). These are
  • $(st^{\prime\prime}-ts^{\prime\prime})(p)[\binom{l}{2}b_{g-\frac{h}{2}-1,g-\frac{l}{2}}\omega_{g-\frac{h}{2}-1}^{(h)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)-\binom{h}{2}b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)]$ if $l\geq 2$ and $h,l$ even;
  • $-(st^{\prime\prime}-ts^{\prime\prime})(p)\binom{h}{2}b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$ if $l=0$ and $h=4k+4$.
@@
$\rho(Q)(\xi_{p}^{2k+3},\xi_{p}^{2k+1})=2\pi i\sum_{\begin{subarray}{c}2\leq l\leq 2k\\ l\text{ even}\end{subarray}}\frac{(2k+1-l)}{l!(4k+4-l)!}(st^{\prime\prime}-ts^{\prime\prime})(p)\cdot\\ \left[\binom{l}{2}b_{g-3-2k+\frac{l}{2},g-\frac{l}{2}}\omega_{g-3-2k+\frac{l}{2}}^{(4k+4-l)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)-\binom{h}{2}b_{g-2-2k+\frac{l}{2},g-\frac{l}{2}-1}\omega_{g-2-2k+\frac{l}{2}}^{(4k+2-l)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)\right]-\\ -2\pi i(st^{\prime\prime}-ts^{\prime\prime})(p)\frac{2k+1}{(4k+4)!}\binom{4k+4}{2}b_{g-2k-2,g-1}\omega_{g-2k-2}^{(4k+2)}(p)\omega_{g-1}(p)$
Notice that every term in the equation is a linear combination of $b_{r,u}$ where $r+u=2g-2k-3$.
Let $\lambda_{k,u}$ be the coefficient of $b_{g-2k-3+u,g-u}$ where $u=1,...,k+1$. We will now see that $\lambda_{k,u}\neq 0$ for every $k,u$. For $1\leq u\leq k$, we have
@@
(55) $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=\sum\limits_{u=1}^{k+1}\lambda_{k,u}b_{g-2k-3+u,g-u}=-\sum\limits_{u=1}^{k+1}\lambda_{k,u}a_{u,2k+3-u}=0$
for every $Q\in Ker(\mu_{2k})$, where $a_{ij}=-b_{g-j,g-i}$ by Remark 4.2. Recall that $Q\in Ker(\mu_{2k})$ if and only if the coefficients $a_{ij}$ satisfy the equations given in Lemma 3.7. Moreover observe that from Theorem 3.1 it follows that we have $k$ linearly independent equations in the variables $a_{u,2k+3-u}$ with $u=1,...,k+1$. We have shown that each coefficient $\lambda_{k,u}$ in equation (55) is different from zero and up to scalar it is the product $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)$. So we can multiply the sections $\omega_{g-u}$ by a non zero scalar multiple, hence equation (55) cannot be linearly dependent from the $k$ equations giving $Ker(\mu_{2k})$. Hence there exists a quadric $Q\in ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$.
@@
Notice that we have shown that the equation (55) gives a hyperplane in $Ker(\mu_{2k})$ that we denote by
@@
Moreover we have just proven that a quadric $Q\in A_{k,0}$ iff satisfies the system of $k+1$ linearly independent equations in $k+1$ variables, given by the equations in Lemma 3.7 defining $Ker(\mu_{2k})$ and equation (55). In particular the coefficients of $Q$ satisfy the equations $b_{g-2k+3+u,g-u}=-a_{u,2k+3-u}=0\quad\forall u=1,...,k+1$.
@@
Let $0\leq k\leq\lfloor\frac{g-3}{2}\rfloor$ and let us now assume $Q\in A_{k,0}$. Then there exist coefficients $\alpha_{k,u}$, all non-zero such that
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=\sum\limits_{u=1}^{k+1}\alpha_{k,u}b_{g-4-2k+u,g-u}\omega_{g-4-2k+u}^{(4k+6-2u)}(p)\omega_{g-u}^{(2u-2)}(p),$
Moreover,
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\Leftrightarrow b_{g-4-2k+u,g-u}=0\quad\forall u=1,...,k+1.$
@@
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=2\pi i[\sum\limits_{u=0}^{2k+2}\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}^{(4k+6-u)}(0)\gamma_{\beta}^{(s)}(0)]\frac{(2k+3-u)}{u!(4k+6-u)!}.$
@@
(56) $\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)$
@@
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{i}^{(m)}\omega_{j}^{(n)})(p)=0\quad\forall m+n\leq 4k+3.$
@@
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=\sum\limits_{u=1}^{k+1}\alpha_{k,u}b_{g-4-2k+u,g-u}\omega_{g-4-2k+u}^{(4k+6-2u)}(p)\omega_{g-u}^{(2u-2)}(p),$
@@
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\Leftrightarrow b_{g-4-2k+u,g-u}=0\quad\forall u=1,...,k+1.$
@@
To show that these are the only ones, we will prove the result when $g$ is even, the case $g$ odd being very similar.